Society of Cleanliness Engineers^{Hitchhiker’s Guide: B Ark}_{Bluesheet}

It’s time to clean up this town!

The unenlightened call you Janitors, but you are Cleanliness Engineers. Without you the world halt, dragged down by its own muck and dust.

Members

• Bob Agrajag (NPC) : A bit paranoid, but he’s great with a mop.